# Supplemetary Material for the paper

### v1

@@ Abstract
Recovering jointly sparse signals in the multiple measurement vectors (MMV) setting is a fundamental problem in machine learning, but traditional methods like MMV orthogonal matching pursuit (M-OMP) and MMV- FOCal Underdetermined System Solver (M-FOCUSS) often require careful parameter tuning or prior knowledge of the sparsity of the signal and/or noise variance. We introduce a novel tuning-free framework that leverages Implicit Regularization (IR) from overparameterization to overcome this limitation. Our approach reparameterizes the estimation matrix into factors that decouple the shared row-support from individual vector entries. We show that the optimization dynamics inherently promote the desired row-sparse structure by applying gradient descent to a standard least-squares objective on these factors. We prove that with a sufficiently small and balanced initialization, the optimization dynamics exhibit a “momentum-like” effect, causing the norms of rows in the true support to grow significantly faster than others. This formally guarantees that the solution trajectory converges towards an idealized row-sparse solution. Additionally, empirical results demonstrate that our approach achieves performance comparable to established methods without requiring any prior information or tuning.
Tuning-Free Structured Sparse Recovery of Multiple Measurement Vectors using Implicit Regularization
Lakshmi Jayalal          Sheetal Kalyani
Indian Institute of Technology, Madras          Indian Institute of Technology, Madras
@@ 1 Introduction
Recovering structured signals from incomplete linear measurements is a central theme in modern signal processing, compressed sensing, and machine learning. A particularly important setting is the multiple measurement vectors (MMV) problem corresponding to signals sharing a common underlying structure, particularly a shared sparse support. This joint sparsity structure, prevalent in applications from neuroimaging (chen2022mmv) to multiple-input multiple-output (MIMO) channel estimation (yang2024mmv) allows for an improved recovery over processing single vectors individually. Classical algorithmic approaches for MMV including MMV subspace pursuit (M-SP) (liu2019mmv), MMV orthogonal matching pursuit (M-OMP) (zhang2022some; chen2020mmv) and MMV- FOCal Underdetermined System Solver (M-FOCUSS)introduced by cotter2005sparse, often depend on the optimal tuning of regularization parameters or require prior knowledge of the sparsity level. These limitations motivate the search for tuning-free methods for achieving row sparsity.
@@
An alternative paradigm based on Implicit Regularization (IR) through overparameterization has recently gained traction (arora2018optimization; arora2019implicit; li2018algorithmic). Specifically, gradient descent is used on an overparameterized or factorized representation of the target signal rather than adding explicit penalty terms (like $\ell_{p}$ norms) to the loss function to encourage sparsity. The dynamics of the optimization process, often combined with specific initialization strategies or early stopping, implicitly guide the solution towards desirable structures like sparsity or low rank. This phenomenon has been analyzed in context of single measurement vector (SMV) sparse recovery by vaskevicius2019implicit; li2021implicit; chou2023more.
@@
Motivated by the success of implicit regularization in the single measurement vector (SMV) domain, this paper explores its application to the row-sparse MMV recovery task. We consider the standard least-squares loss function $\mathcal{L}(\mathbf{X})=\|{\mathbf{Y}-\mathbf{A}\mathbf{X}}\|_{F}^{2}$ and optimize it using gradient descent. We propose a novel overparameterization that leverages the Hadamard product $(\odot)$ and Hadamard power $(^{\odot})$, which represent element-wise multiplication and product, respectively. This reparameterization is $\mathbf{X}=(\mathbf{g}^{\odot 2}\mathbf{1}_{L})\odot\mathbf{V}$, where $\mathbf{1}_{L}=[1,1,\dots,1]_{1\times L}$ is a row vector of ones that replicates the squared elements of $\mathbf{g}\in\mathbb{R}^{N}$ across the columns of $\mathbf{V}\in\mathbb{R}^{N\times L}$. The core idea is that the gradient dynamics acting on the vector $\mathbf{g}$, which captures the shared row support, will implicitly drive most of its elements towards zero. Due to the multiplicative structure of our parameterization, the corresponding rows of $\mathbf{V}$ are simultaneously driven to zero, thereby inducing row sparsity in the reconstructed matrix, $\mathbf{X}$. This parametrization is designed specifically to handle multiple vectors sharing a common structure, distinguishing it from other IR methods that focus on SMV by factorizing a single vector. Furthermore, it contrasts with matrix and tensor factorization methods that induce a low-rank bias rather than the desired row sparsity.
@@
Our main contributions are: (1) we introduce a novel Hadamard parameterization for the MMV problem and derive its gradient descent update rules; (2)we provide a theoretical analysis of the associated gradient flow, establishing formal guarantees that the optimization trajectory converges towards an ideal row-sparse solution; and (3) we further demonstrate the practical effectiveness of our approach through simulations, showing it achieves comparable performance, without any tuning, against established benchmarks like M-OMP, M-SP, and M-FOCUSS.
This research thus aims to bridge the gap between the evolving theory of implicit regularization and its application in the domain of MMV recovery, offering a novel, regularization-free algorithmic framework.
@@ 2 Related Works
This work builds upon research in several related areas, primarily MMV sparse recovery algorithms, the increasingly studied field of IR in machine learning and the techniques used to analyze IR dynamics.
@@
Sparse MMV Recovery Algorithms: The challenge of recovering jointly sparse signals from MMV is addressed by diverse algorithms, many adapted from SMV methods. Approaches range from greedy pursuit algorithms like M-SP, and M-OMP to Bayesian methods like M-FOCUSS. These traditional approaches often rely on explicit sparsity-promoting penalties (like $l_{p}$ norms) or greedy selection rules, sometimes requiring knowledge of the sparsity level or structure. Our work offers an alternative by achieving row sparsity implicitly through optimization dynamics.
@@
Implicit Regularization via Overparameterization: A significant body of recent work focuses on the phenomenon of Implicit Regularization (IR), where the dynamics of gradient descent on an overparameterized model find structured solutions without explicit regularizers (arora2018optimization; arora2019implicit). This effect has been shown to implicitly favor low-rank solutions in matrix sensing (soltanolkotabi2023implicit) and tensor sensing (razin2021implicit; hariz2024implicit). Crucially, implicit regularization has also been investigated for sparse recovery in the SMV setting (li2021implicit; zhao2022high; vaskevicius2019implicit; wu2020implicit). This concept critically extends to achieving structured sparsity, such as group sparsity, through neural reparameterizations, where gradient descent implicitly promotes the desired structure as demonstrated by li2023implicit. The scale of initialization is a pivotal factor in these processes, often determining the nature of the learned solution.However, the factorizations used in prior work are suboptimal for capturing the joint sparsity structure inherent in MMV applications. Our work addresses this gap by introducing a parameterization specifically designed for this context. In this framework, the dynamics of gradient descent on our proposed factors implicitly promote the row-sparse structure critical for joint recovery.
@@
Analysis Techniques for Implicit Regularization: Analyzing the dynamics and convergence of gradient descent in overparameterized models requires specialized techniques. While methods based on Singular Value Decomposition (SVD) (e.g., arora2019implicit; soltanolkotabi2023implicit; li2018algorithmic) and direct element-wise analysis (li2021implicit; zhao2022high; vaskevicius2019implicit; li2023implicit) are effective for some sparse recovery problems, they become intractable for the complex, non-linear parametrization such as the one we propose. We, therefore, adopt a dynamical systems perspective using gradient flow inspired by the works razin2021implicit and hariz2024implicit. This approach characterizes the evolution of the norms of different signal components over time, which reveals a crucial “momentum-like” effect, where components with larger norms grow significantly faster than smaller ones. This effect is the key mechanism driving the implicit row-sparse regularization in our framework.
@@ 3 Preliminaries
We define $\mu$-coherence to characterize the sensing matrix and $\beta$-smoothness to describe the properties of the loss function. We also introduce the unbalancedness and row-unbalancedness constants, which are crucial for analyzing the dynamics of our parameterization.
@@
$\mu$-coherence is a measure of the correlation between the columns of a sensing matrix $\mathbf{A}\in\mathbb{R}^{M\times N}$ in compressed sensing. It is defined as follows:
where $\mathbf{A}_{:i}$ is the $i$-th column of sensing matrix $\mathbf{A}$ and $\langle\cdot,\cdot\rangle$ denotes the inner product.
Throughout this work, we assume that the sensing matrix $\mathbf{A}$ is $\mu$-coherent with $\ell_{2}$-normalized columns.
@@
$\boldsymbol{\beta}$-smoothness. We say that a continuously differentiable function $f(\mathbf{g},\mathbf{V})$ is $\beta$-smooth if its gradient $\nabla f(\mathbf{g},\mathbf{V})$ satisfies
[…]
where $\left\|\begin{matrix}(\mathbf{g},\mathbf{V})-(\tilde{\mathbf{g}},\tilde{\mathbf{V}})\end{matrix}\right\|^{2}=L\|\mathbf{g}-\tilde{\mathbf{g}}\|_{2}^{2}+\|\mathbf{V}-\tilde{\mathbf{V}}\|_{F}^{2}$,
$\left\|\left(\begin{matrix}\nabla_{\mathbf{g}}f(\mathbf{g},\mathbf{V})-\nabla_{\mathbf{g}}f(\tilde{\mathbf{g}},\tilde{\mathbf{V}})\\ \nabla_{\mathbf{V}}f(\mathbf{g},\mathbf{V})-\nabla_{\mathbf{V}}f(\tilde{\mathbf{g}},\tilde{\mathbf{V}})\end{matrix}\right)\right\|^{2}$
$\hskip-80.00012pt=\left\|\nabla_{\mathbf{g}}f(\mathbf{g},\mathbf{V})-\nabla_{\mathbf{g}}f(\tilde{\mathbf{g}},\tilde{\mathbf{V}})\right\|_{2}^{2}$
$\hskip-70.0001pt+\left\|\nabla_{\mathbf{V}}f(\mathbf{g},\mathbf{V})-\nabla_{\mathbf{V}}f(\tilde{\mathbf{g}},\tilde{\mathbf{V}})\right\|_{F}^{2}.$
Definition 3.3.
The constant $\epsilon\geq 0$ be defined as the unbalancedness constant (introduced by razin2021implicit), where $\epsilon=\left|\frac{1}{2}\|\mathbf{g}(t)\|^{2}_{2}-\|\mathbf{V}(t)\|_{F}^{2}\right|$ for any time $t\geq 0$.
Definition 3.4.
The constant $\epsilon_{r}\geq 0$ be defined as the row-unbalancedness constant, where $\epsilon_{r}=\underset{i\in[N]}{\max}\left|\frac{1}{2}{g}_{i}^{2}(t)-\sum\limits_{j}{V}_{ij}^{2}(t)\right|$ for any time $t\geq 0$ and all rows $i$.
With these preliminaries established, we now introduce our IR framework for the MMV problem.
@@ 4 IR-MMV
Consider the task of recovering an MMV $\mathbf{X}\in\mathbb{R}^{N\times L}$. characterized by a common sparsity pattern across its rows. The measurement model is given by:
@@
1:Input: Sensing matrix $\mathbf{A}\in\mathbb{R}^{M\times N}$, measurement matrix $\mathbf{Y}\in\mathbb{R}^{M\times L}$
2:
3:Initialize: $\alpha_{g}=10^{-4},\alpha_{V}=\frac{1}{\sqrt{2L}}\alpha_{g}$, learning rates $\eta_{g}=\eta_{v}=10^{-2}$, number of iterations $T=5\times 10^{6}$, $\mathbf{g}(0)\in\mathbb{R}^{N}$ with $\mathbf{g}(0)=\alpha_{g}\mathbf{1}_{L}$, $\mathbf{V}(0)\in\mathbb{R}^{N\times L}$ with $\mathbf{V}(0)=\alpha_{v}\mathbf{1}_{N\times L}$.
4:Reconstruct initial signal: $\mathbf{X}(0)=(\mathbf{g}(0)^{\odot 2}\mathbf{1}_{L})\odot\mathbf{V}(0)$.
5:
6:for $t=0,1,\dots,T-1$ do
7:  $\boldsymbol{\Lambda}\leftarrow\mathbf{A}^{\top}(\mathbf{Y}-\mathbf{A}\mathbf{X})$
8:  Update parameters:
@@
9:  Reconstruct signal: $\mathbf{X}(t+1)=(\mathbf{g}^{\odot 2}(t+1)\mathbf{1}_{L})\odot\mathbf{V}(t+1)$
10:end for
11:
12:Output: Estimated signal matrix $\mathbf{X}$.
@@
The approach is outlined in Algorithm 1. In line with the analysis of matrix and tensor factorization by gunasekar2017implicit; arora2018optimization; li2018algorithmic; arora2019implicit; razin2021implicit; hariz2024implicit, we also model small learning rate for gradient descent through infinitesimal limit, i.e. through gradient flow:

### v2

# Tuning-Free Structured Sparse Recovery of Multiple Measurement Vectors using Implicit Regularization

@@ Abstract
Recovering jointly sparse signals in the multiple measurement vectors (MMV) setting is a fundamental problem in machine learning, but traditional methods often require careful parameter tuning or prior knowledge of the sparsity of the signal and/or noise variance. We propose a tuning-free framework that leverages implicit regularization (IR) from overparameterization to overcome this limitation. Our approach reparameterizes the estimation matrix into factors that decouple the shared row-support from individual vector entries and applies gradient descent to a standard least-squares objective. We prove that with a sufficiently small and balanced initialization, the optimization dynamics exhibit a “momentum-like” effect where the true support grows significantly faster. Leveraging a Lyapunov-based analysis of the gradient flow, we further establish formal guarantees that the solution trajectory converges towards an idealized row-sparse solution. Empirical results demonstrate that our tuning-free approach achieves performance comparable to optimally tuned established methods. Furthermore, our framework significantly outperforms these baselines in scenarios where accurate priors are unavailable to the baselines.
Machine Learning, ICML, compressed sensing, Implicit regularization, Sparse Signal Processing, Multiple Measurement Vectors, Gradient Flow
@@ 1 Introduction
Recovering structured signals from incomplete linear measurements is a central theme in modern signal processing, compressed sensing, and machine learning. A particularly important setting is the multiple measurement vectors (MMV) problem corresponding to signals sharing a common underlying structure, particularly a shared sparse support. This joint sparsity structure, prevalent in applications from neuroimaging (Chen et al., 2022) to multiple-input multiple-output (MIMO) channel estimation (Yang and Li, 2024) allows for an improved recovery over processing single vectors individually. Classical approaches for MMV including MMV subspace pursuit (M-SP) (Liu et al., 2019), MMV orthogonal matching pursuit (M-OMP) (Zhang et al., 2022; Chen et al., 2020), approximate message passing based MMV (AMP-MMV) (Ziniel and Schniter, 2012) and MMV- FOCal Underdetermined System Solver (M-FOCUSS)introduced by Cotter et al. (2005), often depend on the optimal tuning of regularization parameters or require prior knowledge of the sparsity level. While Bayesian methods like block sparse Bayesian learning extended for MMV (M-BSBL) (Liu et al., 2013) can learn these parameters, they are often susceptible to noise. These limitations motivate the search for robust, tuning-free methods for achieving row sparsity.
@@
An alternative paradigm based on implicit regularization (IR) through overparameterization has emerged with the works of (Arora et al., 2018, 2019; Li et al., 2018). Specifically, gradient descent is used on an overparameterized or factorized representation of the target signal rather than adding explicit penalty terms (like $\ell_{p}$ norms) to the loss function to encourage sparsity. The dynamics of the optimization process, often combined with specific initialization strategies or early stopping, implicitly guide the solution towards desirable structures like sparsity or low rank. This phenomenon has been analyzed in context of single measurement vector (SMV) sparse recovery by Vaskevicius et al. (2019); Li et al. (2021); Chou et al. (2023).
Unlike prior implicit regularization analyses in the SMV setting, such as Vaskevicius et al. (2019); Li et al. (2021); Chou et al. (2023), which factorize a single vector under an $\ell_{2}$ loss, the MMV problem involves recovering a matrix with a shared row-sparse structure. This shared support couples the optimization dynamics across columns and naturally leads to a Frobenius-norm objective, motivating a parameterization tailored to MMV recovery. Building on these observations, we propose an overparameterized formulation for row-sparse MMV recovery and analyze the resulting implicit regularization induced by gradient descent.
@@
We consider the standard least-squares loss function $\mathcal{L}(\mathbf{X})=\|{\mathbf{Y}-\mathbf{A}\mathbf{X}}\|_{F}^{2}$ and optimize it using gradient descent. We propose a novel overparameterization that leverages the Hadamard product $(\odot)$ and Hadamard power $(^{\odot})$, which represent element-wise multiplication and product, respectively. This reparameterization is $\mathbf{X}=(\mathbf{g}^{\odot 2}\mathbf{1}_{L})\odot\mathbf{V}$, where $\mathbf{1}_{L}=[1,1,\dots,1]_{1\times L}$ is a row vector of ones that replicates the squared elements of $\mathbf{g}\in\mathbb{R}^{N}$ across the columns of $\mathbf{V}\in\mathbb{R}^{N\times L}$. The core idea is that the gradient dynamics acting on the vector $\mathbf{g}$, which captures the shared row support, will implicitly drive most of its elements towards zero. Due to the multiplicative structure of our parameterization, the corresponding rows of $\mathbf{V}$ are simultaneously driven to zero, thereby inducing row sparsity in the reconstructed matrix, $\mathbf{X}$. This parametrization is designed specifically to handle multiple vectors sharing a common structure, distinguishing it from other IR methods that focus on SMV by factorizing a single vector. Furthermore, it contrasts with matrix and tensor factorization methods that induce a low-rank bias rather than the desired row sparsity.
@@
Our main contributions are: (1) we introduce a novel Hadamard parameterization for the MMV problem and derive its gradient descent update rules; (2) to the best of our knowledge, we are the first to utilize a Lyapunov-based gradient flow analysis in the context of implicit regularization. This approach departs from standard Grönwall-type arguments, yielding initialization values which are practically feasible. We establish formal guarantees that the trajectory approaches an ideal row-sparse solution under a small and balanced initialization; and (3) we further demonstrate the practical effectiveness of our approach through simulations, showing our approach achieves performance comparable to optimally tuned benchmarks (like M-OMP, M-SP, AMP-MMV, M-BSBL and M-FOCUSS) while overcoming their sensitivity to parameter misspecification.
@@ 2 Related Works
This work builds upon research in several areas, primarily MMV sparse recovery algorithms, IR in machine learning and the techniques used to analyze IR dynamics.
@@
Sparse MMV Recovery Algorithms: The challenge of recovering jointly sparse signals from MMV is addressed by diverse algorithms, many adapted from SMV methods. Approaches range from greedy pursuit algorithms like M-SP, and M-OMP to iteratively reweighted least square based M-FOCUSS and Gaussian approximation approaches like AMP-MMV. These traditional approaches often rely on explicit sparsity-promoting penalties (like $l_{p}$ norms) or greedy selection rules, sometimes requiring knowledge of the sparsity level or structure. Additionally, frameworks like block sparse Bayesian learning extended for MMV (M-BSBL) (Liu et al., 2013) exploit statistical correlations to learn hyperparameters. But, when relying on estimated priors, frameworks like M-BSBL struggle to maintain accurate recovery in low signal to noise ratio (SNR) regimes. Our work offers an alternative by achieving row sparsity implicitly through optimization dynamics.
@@
Implicit Regularization via Overparameterization: A significant body of recent work focuses on the phenomenon of implicit regularization (IR), where the dynamics of gradient descent on an overparameterized model find structured solutions without explicit regularizers (Arora et al., 2018, 2019). This effect has been shown to implicitly favor low-rank solutions in matrix sensing (Soltanolkotabi et al., 2023) and tensor sensing (Razin et al., 2021; Hariz et al., 2024). Crucially, implicit regularization has also been investigated for sparse recovery (Li et al., 2021; Zhao et al., 2022; Vaskevicius et al., 2019; Wu et al., 2020) and structured sparse recovery (Li et al., 2023) in the SMV setting. The scale of initialization is a pivotal factor in these processes, often determining the nature of the learned solution. However, the factorizations used in prior work are suboptimal for capturing the joint sparsity structure inherent in MMV applications. Our work addresses this gap by introducing a parameterization specifically designed for this context. In this framework, the dynamics of gradient descent on our proposed factors implicitly promote the row-sparse structure critical for joint recovery.
@@
Analysis Techniques for Implicit Regularization: Analyzing the dynamics and convergence of gradient descent in overparameterized models requires specialized techniques. While methods based on Singular Value Decomposition (SVD) (e.g., Arora et al. (2019); Soltanolkotabi et al. (2023); Li et al. (2018)) and direct element-wise analysis (Li et al., 2021; Zhao et al., 2022; Vaskevicius et al., 2019; Li et al., 2023) are effective for some sparse recovery problems, they become intractable for the complex, non-linear parametrization such as the one we propose. We, therefore, adopt a dynamical systems perspective using gradient flow inspired by the works Razin et al. (2021) and Hariz et al. (2024). However, unlike these prior studies which rely on Grönwall’s inequality to bound solution trajectories, we employ a Lyapunov-based analysis. This enables us to avoid an exponential dependence on time for initialization and yields practically useful initialization values.
@@ 3 Preliminaries
We begin by establishing the basic notation followed by the formal definitions necessary for our analysis.
With the notation established, we now define $\mu$-coherence to characterize the sensing matrix. We also introduce the unbalancedness and row-unbalancedness constants, which are crucial for analyzing the dynamics of our parameterization.
Definition 3.1 (Mutual coherence).
Let $\mathbf{A}\in\mathbb{R}^{M\times N}$ be a sensing matrix with $\ell_{2}$-normalized columns, i.e., $\|\mathbf{A}_{:i}\|_{2}=1$ for all $i\in[N]$. The mutual coherence of $\mathbf{A}$ is defined as
where $\mathbf{A}_{:i}$ is the $i$-th column of sensing matrix $\mathbf{A}$ and $\langle\cdot,\cdot\rangle$ denotes the inner product. This notion of mutual coherence is standard in compressed sensing and has also been adopted in recent analyses of implicit regularization in overparameterized models, e.g., Li et al. (2021). Throughout this work, we assume that the sensing matrix $\mathbf{A}$ is $\mu$-coherent with $\ell_{2}$-normalized columns.
@@
The constant $\epsilon(t)\geq 0$ be defined as the unbalancedness constant (introduced by Razin et al. (2021)), where $\epsilon(t)=\left|\frac{1}{2}\|\mathbf{g}(t)\|^{2}_{2}-\|\mathbf{V}(t)\|_{F}^{2}\right|$ for any time $t\geq 0$.
@@
The constant $\epsilon_{r}(t)\geq 0$ be defined as the row-unbalancedness constant, where $\epsilon_{r}(t)=\underset{i\in[N]}{\max}\left|\frac{1}{2}{g}_{i}^{2}(t)-\sum\limits_{j}{V}_{ij}^{2}(t)\right|$ for any time $t\geq 0$ and $\forall i\in[N]$.
While defined at any time $t\geq 0$, we show in Lemma 5.1 that both $\epsilon(t)$ and $\epsilon_{r}(t)$ are conserved quantities under the gradient flow dynamics. Consequently, we treat them as time-invariant constants throughout this work, hence dropping the time index. With these preliminaries established, we now introduce our IR framework for the MMV problem.
@@ 4 IR-MMV
Consider the task of recovering an MMV $\mathbf{X}\in\mathbb{R}^{N\times L}$, characterized by a common sparsity pattern across its rows. The measurement model is given by:
@@
Input: Sensing matrix $\mathbf{A}\in\mathbb{R}^{M\times N}$, measurement matrix $\mathbf{Y}\in\mathbb{R}^{M\times L}$
Initialize: $\alpha_{V}=5\times 10^{-4},\alpha_{g}=\alpha_{V}{\sqrt{2L}}$, learning rates $\eta_{g}=\eta_{v}=10^{-4}$, number of iterations $T=5\times 10^{6}$, $\mathbf{g}(0)\in\mathbb{R}^{N}$ with $\mathbf{g}(0)=\alpha_{g}\mathbf{1}_{L}$, $\mathbf{V}(0)\in\mathbb{R}^{N\times L}$ with $\mathbf{V}(0)=\alpha_{v}\mathbf{1}_{N\times L}$.
Reconstruct initial signal: $\mathbf{X}(0)=(\mathbf{g}(0)^{\odot 2}\mathbf{1}_{L})\odot\mathbf{V}(0)$.
for $t=0,1,\dots,T-1$ do
  $\boldsymbol{\Lambda}\leftarrow\mathbf{A}^{\top}(\mathbf{Y}-\mathbf{A}\mathbf{X})$
  Update parameters:
@@
  Reconstruct signal: $\mathbf{X}(t+1)=(\mathbf{g}^{\odot 2}(t+1)\mathbf{1}_{L})\odot\mathbf{V}(t+1)$
end for
Output: Estimated signal matrix $\mathbf{X}$.
@@
The approach is outlined in Algorithm 1. In line with the analysis of matrix and tensor factorization by Gunasekar et al. (2017); Arora et al. (2018); Li et al. (2018); Arora et al. (2019); Razin et al. (2021); Hariz et al. (2024), we also model small learning rate for gradient descent through infinitesimal limit, i.e. through gradient flow: